# Quantitative agreement between experiment and theory for Vibrational Circular Dichroism enhanced by electronically excited states

### v1

@@ 5.4 TDDFT calculations of the enhancement
When we restrict the optimization to just the two lowest excited states, the agreement with experiment remains strong: the overall simVCD falls to 0.54 (versus 0.68 when all three states are adjusted), yet the spectral profile is still well reproduced. Most of the discrepancy arises in the 1250–1400 cm-1 region, and in a more pronounced negative peak around 1100 cm-1. The right-hand panel of  Figure 9 demonstrate contributions of each excited state, revealing that the same VCD band can flip sign depending on the state. For instance, the first excited state produces a negative feature near 1300 cm-1, while both the second and third excited states give positive peaks at that frequency. Moreover, different states dominate distinct spectral regions: in the 1500–1200 cm-1 region a superposition of the first and second excited state contributions reproduces the overall band shape, the 1200-1100 cm-1 region is well described by the second state, and finally 1100-1000 cm-1 is closer to the shape of the third state contribution.
For Ni(sp)Cl2, the TDDFT results using the BHandH functional are essentially identical to those obtained with SA-CASSCF: only the first excited state contributes to the enhancement, giving a final simVCD of 0.45 (see  Figure 11) with optimized energy 0.23 eV. The non-adiabatic couplings computed with BHandH closely match those from SA-CASSCF for Ni(sp)Cl2, which explains the close result ( Figure 12). We again observe a large negative peak at 1300 cm-1, similar to SA-CASSCF case.
@@
(a) NACs between the ground and first excited state
(b) NACs between the ground and second excited state
(c) NACs between the ground and third excited state
@@ 5.5 Assignment of absolute configuration
An critical test of our semi-empirical approach is its ability to assign absolute configuration reliably. To probe this, we inverted the experimental VCD spectrum and repeated the energy optimization procedure. For Co(sp)Cl2 the optimization drives the first excited-state energy down to an unphysically low 0.08 eV, where perturbation theory breaks down (see Figure 13). The other two states converge to 0.37 eV and contribute negligibly. The resulting VCD intensities are dramatically overestimated, especially in the 1100–1000 cm-1 region, which is a clear evidence of a breakdown in the fitted enantiomer spectrum. For Ni(sp)Cl2 the optimized simVCD remains moderately high (0.32) and the spectral shape is closer to the inverted spectrum than in the Co case. However, state analysis reveals that only the third excited state contributes constructively to the fit. Moreover, optimization drives the energy of this state to a low energy of 0.18 eV, while the first two states end up with degenerate energies of 0.37 eV. In general, we show that fitting the spectrum of another enantiomer with the proposed optimization procedure leads to results, which can be readily identified as being erroneous. This makes it possible to rely on the tuned enhanced calculation result for robust and unambiguous assignment of absolute configuration, provided that the set of low-lying states used in the tuning is carefully selected and validated.
@@
Figure 13: Enhanced VCD spectra for the enantiomer of Co(sp)Cl2 (left). The calculated spectrum is obtained using SA-CASSCF with three excited states and optimized excitation energies (red); the enantiomeric VCD spectrum is obtained by inversion (black). Individual contributions of each state are given on the right. The second and the third state have negligible contributions with simVCD -0.05 (red) and 0.02 (purple), respectively.
@@
Figure 14: Enhanced VCD spectra for the enantiomer of Ni(sp)Cl2 (left). The calculated spectrum is obtained using SA-CASSCF states and optimized excitation energies (red); the enantiomeric VCD spectrum is obtain by inversion (black). Individual contributions of each state are given on the right. The first state has a negative similarity -0.12 (dark red); the overall positive similarity is coming from the third state (purple, simVCD = 0.34)
5.6 Understanding the mechanism of enhancement
@@
Another advantage of the formalism presented in this work is the possibility to understand the mechanism of the enhancement in full detail. To this end we note that we can distinguish enhancement due to the magnetic and electric dipole components of the electronic transition moments:
@@
We observe for all the calculated enhanced spectra that the electric dipole contribution is negligible. This arises from the small electric dipole transition moments associated with pure $d-d$ transitions. The decomposition of the enhanced spectra into contributions for Co(sp)Cl2 and Ni(sp)Cl2 is demonstrated in Figure 15. We clearly observe the magnetic dipole enhancement dominating the signal, making the accuracy of the APT tensor (from the MFP calculation) and the magnetic dipole transition moments between electronic states more critical than the AAT tensor and the electric dipole transition moments.
@@ 6 Conclusions
In this work, we have applied Nafie’s sum-over-states approach to calculate the VCD enhancement in transition metal complexes. We studied rigid sparteine complexes as a proof-of-concept application of vibronic coupling theory. The next step will be to extend this approach to flexible complexes, where conformational variability introduces additional complications.
We find that the enhancement is extremely sensitive to the excitation energies, making it challenging for quantum chemistry methods to predict these values with sufficient accuracy. Rather than pushing for even more accurate excited state calculations, we propose treating these key excitation energies as adjustable parameters. By optimizing just a few values, we can reproduce experimental VCD spectra with high similarity, achieving simVCD over 0.4 for both SA-CASSCF and TDDFT.
@@
Our results suggest that TDDFT can perform comparably to CASSCF, provided that an appropriate functional is chosen. In our study, BHandH delivers the best VCD enhancements, likely due to its high fraction of Hartree–Fock exchange, which is crucial for accurately describing $d-d$ transitions in transition metal complexes. In fact, for Ni(sp)Cl2, the non-adiabatic couplings calculated with BHandH closely match those of SA-CASSCF, leading to nearly identical spectra with SA-CASSCF states. For Co(sp)Cl2, we observe some discrepancies in the couplings to the first and third excited states between the two methods, accounting for the slight differences in their respective contributions.
An important observation, noted in earlier studies, is the local character of the VCD enhancement. Here, we show that significant nonadiabatic couplings occur only between the transition metal and its close neighbors: two nitrogens and two chlorines in the sparteine case. Consequently, we assume that for larger systems, one could compute these couplings on a truncated model without substantial loss of accuracy. We plan to explore this idea in future work.
Our findings furthermore demonstrate that problems to computationally reproduce enhanced VCD spectra were due to inaccuracies in just a few key parameters. Making these parameters tunable leads to an efficient and practical scheme to determine absolute configurations by means of enhanced VCD spectra in transition metal systems.
@@ 7 Acknowledgements
We acknowledge financial support from NWO in the framework of the Fund New Chemical Innovations (NWO Project Nr.731.014.209) and the National Grow Fund Quantum Delta NL (NWO Project Nr. NGF.1582.22.036). SAJ acknowledges support of the Chemistry & Chemical Engineering Research Center of Iran (CCERCI) in the framework of a student exchange program. We furthermore acknowledge NWO for providing access to Snellius, hosted by SURF through the Computing Time on National Computer Facilities call for proposals.

### v2

@@ 5.4 TDDFT calculations of enhancements
When we restrict the optimization to just the two lowest excited states, the agreement with the experiment remains very good. The simVCD value is reduced to 0.54 compared to 0.67 when all three states are adjusted – which is still quite acceptable in terms of reliability of the assignment of the absolute configuration – but equally important is that the overall spectral profile is still well reproduced. Most of the discrepancy arises in the 1250–1400 cm-1 region, and from the more pronounced negative peak around 1100 cm-1. The righthand panel of  Figure 9 shows the contributions of each excited state. Comparison of these contributions emphasizes once again the observation made previously when discussing the SA-CASSCF results that each state has a unique enhancement fingerprint both in sign and magnitude. For instance, the first excited state produces a negative feature near 1300 cm-1, while both the second and third excited states give positive peaks at that frequency. Moreover, different states dominate distinct spectral regions: in the 1500–1200 cm-1 region a superposition of the first and second excited state contributions reproduces the overall band shape, the 1200-1100 cm-1 region is well described by the second state, and finally 1100-1000 cm-1 is closer to the shape of the third state contribution.
For Ni(sp)Cl2, the TDDFT results using the BHandH functional are essentially identical to those obtained with SA-CASSCF: only the first excited state contributes to the enhancement, giving a final simVCD value of 0.45 (see  Figure 11) and an optimized energy of 0.23 eV. The NACs computed with BHandH closely match those from SA-CASSCF for Ni(sp)Cl2Figure 12), which explains that both approaches lead to the same result and confirms our explanation for the differences found for the Co complex. Similar to what was found with SA-CASSCF, we also observe with TDDFT a large negative peak at 1300 cm-1.
@@
(a) NACs between ground and first excited state
(b) NACs between ground and second excited state
(c) NACs between ground and third excited state
@@ 5.5 Assignment of absolute configuration
A critical test of the applicability of our approach is its ability to come to a reliable assignment of the absolute configuration. In order to explore this important aspect, we inverted the experimental VCD spectrum and repeated the energy optimization procedure using our SA-CASSCF results. For Co(sp)Cl2 the results of such an optimization are shown in Figure 13. The first and foremost conclusion that can be drawn is that such an optimization leads to a simVCD value of 0.17 which is (i) clearly much lower than obtained for the correct configuration (0.69) and (ii) too low to come to a reliable assignment of the absolute configuration. Looking into more detail into the results of the optimization, we find that it leads to an nonphysically low value of the excitation energy of the first excited state of 0.08 eV, a value for which perturbation theory would in any case no longer be valid. Furthermore, it is only this state that contributes to the enhancements; the excitation energies of the other two states converge to 0.37 eV and these states contribute negligibly. The VCD spectrum that then comes out is a spectrum in which band intensities are dramatically overestimated, especially in the 1100–1000 cm-1 region, which is a further indication that it is not possible to fit the observed experimental spectrum with the calculated input of the opposite enantiomer.
For Ni(sp)Cl2 (Figure 14) energy optimization similarly leads to a smaller simVCD value though less pronounced than observed for the Co complex (0.32 versus 0.51). We also find that the overall shape of the fitted spectrum does not deviate as much as found for the Co complex, although there are regions where there are significant discrepancies with the experimental spectrum, especially in 1300-1100 cm-1 range. Further support for the conclusion that the observed spectrum cannot be attributed to the enantiomer – as one would conclude if one would not know the absolute configuration of the complex – is the observation that optimization reduces the excitation energy of the third excited state to a low value of 0.18 eV while the first two states end up with a similar energy of 0.37 eV. Such an inversion would be at odds with a priori expectations on the ordering of these states. A consequence of these state energies is that only the third excited state contributes constructively to the fit.
We thus conclude that fitting the spectrum of the opposite enantiomer with the proposed optimization procedure leads to results that can readily be identified as being inconsistent with the experiment. This makes it possible to rely on the results of the tuned enhanced calculation to come to a robust and unambiguous assignment of the absolute configuration, provided that the set of low-lying states used in the tuning is carefully selected and validated.
@@
Figure 13: Left panel: Enhanced VCD spectra for the enantiomer of Co(sp)Cl2 (left). The calculated spectrum is obtained using SA-CASSCF with three excited states and optimized excitation energies (red); the enantiomeric VCD spectrum is obtained by inversion of the original spectrum (black). Right panel: Individual contributions of each state. The second and third state have negligible contributions and give rise to spectra with a simVCD value of -0.05 (red) and 0.02 (purple), respectively.
@@
Figure 14: Left panel: Enhanced VCD spectra for the enantiomer of Ni(sp)Cl2 (left). The calculated spectrum is obtained using SA-CASSCF states and optimized excitation energies (red); the enantiomeric VCD spectrum is obtain by inversion of the original spectrum (black). Right panel: Individual contributions of each state. The first state gives rise to a spectrum with a simVCD value of -0.12 (dark red); the overall positive simVCD value is primarily originating from the third state (purple, simVCD = 0.34)
5.6 Understanding the enhancement mechanism
@@
We conclude our analysis by pointing out that another advantage of the formalism presented in this work is that it offers the possibility to come to a more detailed understanding of the enhancement mechanism. To this end we notice that we can distinguish enhancement due to the magnetic and electric dipole components of the electronic transition moments:
@@
For the case at hand we find for all the calculated enhanced spectra that the electric dipole contribution is negligible. This is as expected since electric dipole transition moments associated with pure $d$-$d$ transitions are formally forbidden (Laporte selection rule). The decomposition of the enhanced spectra into contributions for Co(sp)Cl2 and Ni(sp)Cl2 is demonstrated in Figure 15. We clearly observe the magnetic dipole enhancement dominating the signal, making the accuracy of the APT tensor (from the MFP calculation) and the magnetic dipole transition moments between electronic states more critical than the AAT tensor and the electric dipole transition moments. Although for the present case such a result is as expected, for a more general case such a decomposition will clearly contribute to a further elucidation of the factors that contribute to an observed enhancement.
@@ 6 Conclusions
In this work, we have applied the sum-over-states approach introduced by Nafie to calculate enhancements of band in VCD spectra of transition metal complexes. We find that these enhancements are extremely sensitive to the excitation energies, making it challenging for quantum chemistry methods to predict these values with sufficient accuracy. Rather than pushing for even more accurate excited state calculations, we propose treating these key excitation energies as adjustable parameters. By optimizing just a few values, we can reproduce experimental VCD spectra with a high similarity, achieving simVCD values over 0.4 for both SA-CASSCF and TDDFT.
@@
Our results indicate that TDDFT can perform comparably to CASSCF, provided that an appropriate functional is chosen. In our study, BHandH delivers the best VCD enhancements, likely due to its high fraction of Hartree–Fock exchange, which is crucial for accurately describing $d$-$d$ transitions in transition metal complexes. In fact, for Ni(sp)Cl2, the non-adiabatic couplings calculated with BHandH closely match those of SA-CASSCF, leading to nearly identical spectra with SA-CASSCF states. For Co(sp)Cl2, we observe some differences in the couplings to the first and third excited states between the two methods, accounting for the slight differences in their respective contributions.
An important observation noted in earlier studies as well is the local character of the VCD enhancement. Here, we show that significant non-adiabatic couplings occur only between the transition metal and its close neighbors: two nitrogens and two chlorines in the sparteine case. Consequently, we assume that for larger systems, one could compute these couplings on a truncated model without substantial loss of accuracy. We plan to explore this idea in future work. A further direction we are working on concerns the fact that in the present study we have studied rigid sparteine complexes as a proof-of-concept application of vibronic coupling theory. The next step will be to extend this approach to flexible complexes, where conformational variability introduces additional complications.
@@ 7 Acknowledgements
We acknowledge financial support from NWO in the framework of the Fund New Chemical Innovations (NWO Project Nr.731.014.209) and the National Grow Fund Quantum Delta NL (NWO Project Nr. NGF.1582.22.036). We furthermore acknowledge NWO for providing access to Snellius, hosted by SURF through the Computing Time on National Computer Facilities call for proposals.